OpenStudy (anonymous):

I have to solve the system 7x - 3y = 20 5x + 3y = 16 I solved it once; 7x - 3y = 20 + 5x + 3y = 16 = 12x = 36 = x = 3 7(3) = 3y = 20 = 21 - 3y = 20 21 - 21 - 3y = 21 - 21 = 3y = -1 = y = 1/3 so (3, 1/3) But I cant figure out how to solve it for a nother way and get the same answer.

OpenStudy (anonymous):

Example: 2X+6Y=10 3X-2Y=12 There is another method called substitution, in which you manipulate one of the equation to isolate a variable. For example 2X+6Y=10 can be manipulated into X=5-3Y Then you substitute the second part of the equation, 5-3Y in the second equation, 3X-2Y=12, resulting in 3(5-3Y)-2Y=12. Then you can use simple algebra to figure out X and Y

OpenStudy (anonymous):

I know about substitution, and I have attempted it several times in several different manners but I cannot come up with the same answer as the first.

OpenStudy (anonymous):

If at all.

hero (hero):

There's at least five different ways of solving these 1. Graphing 2. Substitution 3. Elimination 4. Matrices 5. Cramer's Method Pick one

hero (hero):

There's also slope-intercept method as well